# Reinforced MLLM: A Survey on RL-Based Reasoning in Multimodal Large Language Models

### v1

@@ 2.3.2. Methodologies in MLLM Reasoning
Inspired by the significant potential of test-time scaling in the field of NLP for solving complex language tasks (Snell et al., 2024; Guo et al., 2025; Jaech et al., 2024), recent studies have sought to enhance the reasoning capabilities of MLLMs. Broadly, current MLLM reasoning methods fall into two main categories:
@@
The first method focuses on guiding models to generate high-quality CoT data that adheres to specific steps. This can be achieved either through structured reasoning templates—such as summaries, descriptions, logical reasoning, and conclusions—or by leveraging advanced search algorithms. Notable examples include CoT prompting techniques (Zhao et al., 2023; Fei et al., 2024; Chen et al., 2023b), planning-based approaches like Best-of-N (Chen et al., 2024c; Feng et al., 2023; Xu et al., 2024; Wang et al., 2022), Monte Carlo Tree Search (MCTS) (Yao et al., 2024; Tian et al., 2024; Ding et al., 2023; Xie et al., 2024), and Beam Search (Wu and Xie, 2024; Yao et al., 2023; Xie et al., 2023). The generated CoT data is then utilized for SFT, enabling MLLMs to emulate high-quality, structured reasoning during inference. While this approach ensures the quality of CoT data and enables effective integration of multi-modal information, it suffers from two key limitations: the labor-intensive nature of CoT annotation and the risk of catastrophic forgetting during the fine-tuning process.
@@
The second category builds on the DeepSeek R1 (Guo et al., 2025) reinforcement learning paradigm, which integrates large-scale RL with structured and outcome-driven reward functions. This framework enables large language models to automatically generate human-like reasoning processes with remarkable complexity (Team et al., 2025; Liu et al., 2025d; Zhao et al., 2025; Meng et al., 2025; Peng et al., 2025), eliminating the need for labor-intensive CoT annotations. Additionally, it demonstrates strong generalization capabilities when handling out-of-distribution (OOD) data.However, current RL-based reasoning methods for MLLMs are primarily adapted from NLP, with the generated reasoning trajectories predominantly presented in textual form, neglecting interactions among diverse modalities such as images, videos, and audio in MM-CoT. Additionally, the lack of effective reasoning path rewards in current RL-based reasoning often leads to CoT outputs containing excessive redundant tokens.
@@ 2.3.3. SFT Memory, RL Generalization
SFT-based reasoning methods activate the model’s inherent reasoning capabilities by identifying precise cognitive templates and directly constructing high-quality reasoning trajectories. In contrast, RL-based approaches focus on leveraging RL to search for optimal reasoning paths among numerous possibilities, treating the cultivation of reasoning abilities as a process that requires extensive training to explore effective reasoning patterns. This distinction is now widely recognized by the principle: SFT for memory, RL for generalization (Chu et al., 2025). SFT offers strong inductive biases and precise knowledge injection, whereas RL unleashes the model’s potential for abstract, adaptive, and robust reasoning under uncertainty. In the context of MLLMs, where cross-modal alignment and dynamic understanding are critical, this dichotomy becomes even more pronounced. Therefore, advancing RL-based reasoning adaptation techniques tailored for multimodal scenarios is essential. This study contributes to this direction by providing a comprehensive review and analysis of RL-based reasoning strategies in MLLMs.
@@ 3. Key Design and Optimization of RL Algorithms in LLMs/MLLMs
RL methods in LLMs primarily include value-based (e.g. PPO (Schulman et al., 2017)) and value-free (e.g. GRPO (Shao et al., 2024) ) approaches. While GRPO exhibits entropy collapse and instability, PPO faces challenges in long-chain reasoning. Recent optimizations address these limitations, demonstrating improved performance in LLMs, with potential applicability to MLLMs for enhanced training efficiency and stability.
3.1. Value-Free
Value-free reinforcement learning methods, such as GRPO (Shao et al., 2024), have demonstrated remarkable effectiveness in LLM training by eliminating the need for value model computation. Instead, these approaches derive advantages from trajectory-level rewards uniformly assigned to each token, providing stability through group-based reward averaging. This characteristic proves particularly valuable in scenarios where training value models is challenging, making value-free methods highly successful in complex tasks like long-chain reasoning (long-CoT). While these methods offer implementation simplicity and robustness, they also reveal fundamental limitations, such as entropy collapse and reward noise, which recent advancements aim to address through novel optimization techniques.
@@ 3.1.1. DAPO
Compared to GRPO, DAPO (Yu et al., 2025) significantly improves the model’s generation diversity and training efficiency by introducing an asymmetric clipping strategy, a dynamic sampling mechanism, token-level policy gradient loss, and overlong reward shaping. The optimization formula is as follows:
(7) $J_{\mathrm{DAPO}}(\theta)$ $=\mathbb{E}_{(q,a)\sim D,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}{(\cdot|q)}}$
$\left[\frac{1}{\sum_{i=1}^{G}\|o_{i}\|}\sum_{i=1}^{G}\sum_{t=1}^{\|o_{i}\|}\min\left(r_{i,t}(\theta)\hat{A}_{i,t},\operatorname{clip}(r_{i,t}(\theta),1-\epsilon_{\mathrm{low}},1+\epsilon_{\mathrm{high}})\hat{A}_{i,t}\right)\right],$
where $\hat{A}_{i,t}=\frac{r_{i}-\max(\{R_{i}\}_{i=1}^{G})}{\mathrm{std}(\{R_{i}\}_{i=1}^{G})}$, and the other symbols are consistent with those in Equation 5.
@@
In GRPO, the clipping range is typically symmetric, i.e., $\epsilon_{\mathrm{low}}=\epsilon_{\mathrm{high}}$. However, this symmetric clipping strategy limits the exploration of low-probability tokens, potentially leading to entropy collapse. To address this issue, DAPO introduces the Clip-Higher strategy in Equation 7, decoupling the clipping bounds into $\epsilon_{\mathrm{low}}$ and $\epsilon_{\mathrm{high}}$. This asymmetric clipping strategy significantly enhances the exploration of low-probability tokens while avoiding excessive compression of high-probability tokens.
@@
To address the issue where gradient signals vanish if all samples in a batch have accuracy rates of either 1 or 0, DAPO proposes a dynamic sampling strategy. By filtering out samples with accuracy rates of 0 or 1, this strategy ensures that every batch contains samples with valid gradient signals, represented as: $0<|\{o_{i}|\text{is equivalent}(a,o_{i})\}|<G.$ The authors demonstrate that dynamic sampling not only improves training efficiency but also reduces the number of steps required for model convergence.
@@
GRPO computes loss at the sample level, averaging losses across tokens within each sample before aggregating across different samples. This approach poses challenges in long-chain reasoning scenarios, as the contribution of tokens in longer sequences to the overall loss is diluted, making it difficult for the model to learn high-quality reasoning patterns. To address this, DAPO proposes token-level policy gradient loss, which ensures that tokens in longer sequences contribute gradients more fairly, thereby avoiding excessive penalties on low-quality long samples.
@@
To alleviate the issue of ”overthinking”, Overlong Reward Shaping strategy implements a progressive length-dependent penalty system that adjusts rewards for truncated outputs while discouraging excessive verbosity through scaled punishments proportional to response length beyond the threshold, expressed as follows:
@@ 3.1.3. CPPO
While the completion pruning strategy significantly reduces computational overhead, it also introduces the issue of underutilization of GPU resources. After pruning, the reduced number of retained completions fails to fully leverage the parallel computing capabilities of GPUs. To address this, CPPO combines the remaining pruned completions with high-quality completions from new queries, filling the GPU processing pipeline. This ensures that each GPU device operates at full capacity, thereby maximizing the utilization of GPU parallel computing power.
3.2. Value-Based
Value-based methods, exemplified by PPO, excel through precise, step-wise credit assignment—a crucial feature for complex reasoning tasks where localized errors can cascade. Unlike Monte Carlo-based estimation in value-free approaches, value-based methods provide lower-variance token-level estimates and achieve better sample efficiency via generalization. Although these methods face significant challenges in long chain-of-thought reasoning tasks, recent research has introduced innovative optimization techniques to enhance both training stability and performance. When successfully trained, value-based approaches promise higher performance ceilings, presenting a compelling trade-off in reinforcement learning for LLMs and offering promising potential for transfer to multimodal scenarios.
3.2.1. Open-Reasoner-Zero
Open-Reasoner-Zero (ORZ) (Hu et al., 2025) demonstrates through extensive experiments that by using the vanilla PPO algorithm combined with GAE ($\lambda=1$, $\gamma=1$) and a simple rule-based reward function, significant improvements in response length and benchmark performance can be achieved without any KL regularization. Additionally, the authors highlight the importance of increasing both the quantity and diversity of training data for enhancing model performance. By utilizing a carefully curated large-scale and diverse dataset, the model exhibits consistent scaling capabilities on both training and test sets, with no signs of saturation. Compared to DeepSeek-R1-Zero, ORZ achieves comparable performance with only one-tenth of the training steps.
@@ 3.2.2. VC-PPO
To enable traditional reinforcement learning algorithms like PPO to handle long Chain-of-Thought (CoT) tasks, VC-PPO (Yuan et al., 2025) introduces optimizations in two key areas: value initialization bias and Decoupled-GAE.
@@
The value function $V(s)$ in PPO is typically initialized using a reward model during the early stages of training. However, this initialization method can lead to significant bias, especially in long-sequence tasks. In the initial phase, since the reward model tends to underestimate the scores of earlier tokens, the advantage function $A_{t}$ exhibits a positive bias when estimating the contributions of these tokens. This bias encourages the policy to favor shorter outputs, thereby undermining the model’s ability to generate long CoT sequences. To address this issue, VC-PPO proposes Value Pretraining. By offline training the value model under a fixed policy and using Generalized Advantage Estimation (GAE) with $\lambda=1.0$ to ensure unbiased returns, the value model achieves a more accurate estimation of expected rewards.
@@ 4. RL-Based MLLM Reasoning
KIMI K1.5 (Team et al., 2025) KIMI K1.5 SFT&RL OPMD Univ. Reason T
@@
Visual-RFT (Liu et al., 2025d) Qwen2-VL-2B RL GRPO Detection&CLS T
VisualThinker-R1-Zero (Zhou et al., 2025) Qwen2-VL-2B RL GRPO Spatial Reason T
@@
Seg-Zero (Liu et al., 2025c) Qwen2.5-VL-3B RL GRPO Segmentation T
@@
MM-Eureka-8B (Meng et al., 2025) InternVL2.5-Instruct-8B RL RLOO Math T
MM-Eureka-Zero-38B (Meng et al., 2025) InternVL2.5-Instruct-38B RL RLOO Math T
@@
R1-VL (Zhang et al., 2025) Qwen2.5-VL-7B SFT&RL StepGRPO Univ. Reason T&I Desc
@@
OThink-MR1 (Liu et al., 2025e) Qwen2-VL-2/7B-Instruct RL GRPO-D Univ. Reason T
OpenVLThinker-7B (Deng et al., 2025a) Qwen2.5-VL-7B-Instruct SFT&RL GRPO Math T
MetaSpatial (Pan and Liu, 2025) Qwen-VL 3B/7B RL GRPO Univ. Reason T
UI-R1 (Lu et al., 2025) Qwen2.5-VL-3B RL GRPO Action Predict T
@@
Q-Insight (Li et al., 2025b) Qwen-2.5-VL-7B RL GRPO Image Quality T&I Desc
@@
Qwen2-Audio-Instruct-RL (Li et al., 2025a) Qwen2-Audio-7B-Instruct RL GRPO Audio QA T

### v2

@@ 2.3.2. Methodologies in MLLM Reasoning
Inspired by the significant potential of reasoning in NLP for solving complex language tasks (Guo et al., 2025; Jaech et al., 2024), recent studies have sought to enhance the reasoning capabilities of MLLMs. Broadly, current MLLM reasoning methods fall into two main categories:
@@
The first method focuses on guiding models to generate high-quality CoT data that adheres to specific steps. This can be achieved either through structured reasoning templates—such as summaries, descriptions, logical reasoning, and conclusions—or by leveraging advanced search algorithms. Notable examples include CoT prompting techniques (Zhao et al., 2023; Fei et al., 2024), planning-based approaches like Best-of-N (Chen et al., 2024b; Xu et al., 2024), Monte Carlo Tree Search (MCTS) (Yao et al., 2024; Xie et al., 2024), and Beam Search (Wu and Xie, 2024; Yao et al., 2023). The generated CoT data is then utilized for SFT, enabling MLLMs to emulate high-quality, structured reasoning. While this approach ensures the quality of CoT data and enables effective integration of multi-modal information, it suffers from two key limitations: the labor-intensive nature of CoT annotation and the risk of catastrophic forgetting during fine-tuning.
@@
The second category builds on the DeepSeek R1 (Guo et al., 2025) reinforcement learning paradigm, which integrates large-scale RL with structured and outcome-driven reward functions. This framework enables large language models to automatically generate human-like reasoning processes with remarkable complexity (Team et al., 2025a; Liu et al., 2025f; Zhao et al., 2025b; Meng et al., 2025; Peng et al., 2025), eliminating the need for labor-intensive CoT annotations. Additionally, it demonstrates strong generalization capabilities when handling OOD data.However, current RL-based reasoning methods for MLLMs are primarily adapted from NLP, with the generated reasoning trajectories predominantly presented in textual form, neglecting interactions among diverse modalities such as images, videos, and audio in MM-CoT. Additionally, the lack of effective reasoning path rewards in current RL-based reasoning often leads to CoT outputs containing excessive redundant tokens.
@@ 2.3.3. SFT Memory, RL Generalization
SFT-based reasoning methods activate the model’s inherent reasoning capabilities by identifying precise cognitive templates and directly constructing high-quality reasoning trajectories. In contrast, RL-based approaches focus on leveraging RL to search for optimal reasoning paths among numerous possibilities, treating the cultivation of reasoning abilities as a process that requires extensive training to explore effective reasoning patterns. This distinction is now widely recognized by the principle: SFT for memory, RL for generalization (Chu et al., 2025b). SFT offers strong inductive biases and precise knowledge injection, whereas RL unleashes the model’s potential for abstract, adaptive, and robust reasoning under uncertainty. In the context of MLLMs, where cross-modal alignment and dynamic understanding are critical, this dichotomy becomes even more pronounced. Therefore, advancing RL-based reasoning adaptation techniques tailored for multimodal scenarios is essential. This study contributes to this direction by providing a comprehensive review and analysis of RL-based reasoning strategies in MLLMs.
@@ 3. Key Design and Optimization of RL Algorithms in LLMs/MLLMs
In RL for LLMs, methods can be classified into value-model-free such as GRPO (Shao et al., 2024) and value-model-based such as PPO (Schulman et al., 2017) approaches, depending on whether a value model is learned (Yuan et al., 2025a). While GRPO exhibits entropy collapse and instability, PPO faces challenges in long-chain reasoning. Recent optimizations address these limitations, demonstrating improved performance in LLMs, with potential applicability to MLLMs for enhanced training efficiency and stability.
3.1. Value-Model-Free
Value-model-free RL methods, such as GRPO (Shao et al., 2024), have demonstrated remarkable effectiveness in LLM training by eliminating the need for value model computation. Instead, these approaches derive advantages from trajectory-level rewards uniformly assigned to each token, providing stability through group-based reward averaging. This characteristic proves particularly valuable in scenarios where training value models is challenging, making value-model-free methods highly successful in complex tasks like long-chain reasoning. While these methods offer implementation simplicity and robustness, they also reveal fundamental limitations, such as entropy collapse and reward noise, which recent advancements aim to address through novel optimization techniques.
@@ 3.1.1. DAPO
Dynamic sAmpling Policy Optimization (DAPO) (Yu et al., 2025b) significantly improves the model’s generation diversity and RL training efficiency by introducing an asymmetric clipping strategy, a dynamic sampling mechanism, a token-level policy gradient loss, and extended reward shaping. The optimization formula is as follows:
$J_{\mathrm{DAPO}}(\theta)$ $=\mathbb{E}_{(q,a)\sim D,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}{(\cdot|q)}}$
(7) $\left[\frac{1}{\sum_{i=1}^{G}\|o_{i}\|}\sum_{i=1}^{G}\sum_{t=1}^{\|o_{i}\|}\min\left(r_{i,t}(\theta)\hat{A}_{i,t},\operatorname{clip}(r_{i,t}(\theta),1-\epsilon_{\mathrm{low}},1+\epsilon_{\mathrm{high}})\hat{A}_{i,t}\right)\right]$
$\mathrm{s.t.}\quad$ $0<\left|\{o_{i}\mid\mathtt{is\_equivalent}(a,o_{i})\}\right|<G,$
where $\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(\{R_{i}\}_{i=1}^{G})}{\mathrm{std}(\{R_{i}\}_{i=1}^{G})}$, and the other symbols are consistent with those in Equation 5.
@@
In GRPO, the clipping range is typically symmetric, i.e., $\epsilon_{\mathrm{low}}=\epsilon_{\mathrm{high}}$. However, this symmetric clipping strategy limits the exploration of low-probability tokens, potentially leading to entropy collapse. To address this issue, DAPO introduces the Clip-Higher strategy in Equation 3.1.1, decoupling the clipping bounds into $\epsilon_{\mathrm{low}}$ and $\epsilon_{\mathrm{high}}$. This asymmetric clipping strategy significantly increases the value of $\epsilon_{\mathrm{high}}$, thereby allowing for greater flexibility in enhancing low-probability tokens.
@@
To address the issue where gradient signals vanish if all samples in a batch have accuracy rates of either 1 or 0, DAPO proposes a dynamic sampling strategy. By filtering out samples with accuracy rates of 0 or 1, this strategy ensures that every batch contains samples with valid gradient signals, represented as: $0<|\{o_{i}|\text{is equivalent}(a,o_{i})\}|<G.$ The authors demonstrate that dynamic sampling can substantially improve training efficiency.
@@
GRPO computes loss at the sample level, averaging losses across tokens within each sample before aggregating across different samples. This approach poses challenges in long-chain reasoning scenarios, as the contribution of tokens in longer sequences to the overall loss is diluted, making it difficult for the model to learn high-quality reasoning patterns. To address this, DAPO proposes token-level policy gradient loss, which ensures that tokens in longer sequences contribute gradients more fairly.
@@
To address reward noise caused by improper shaping of truncated samples, the Overlong Reward Shaping strategy introduces a progressive length-dependent penalty mechanism. This adjusts rewards for truncated outputs and discourages excessive verbosity by applying scaled punishments proportional to response lengths beyond a threshold, as follows:
@@ 3.1.3. CPPO
While the completion pruning strategy reduces computational overhead, it causes underutilization of GPU resources. After pruning, the reduced number of retained completions fails to fully leverage the parallel computing capabilities of GPUs. To address this, CPPO combines the remaining pruned completions with high-quality completions from new queries, filling the GPU processing pipeline. This ensures that each GPU device operates at full capacity, thereby maximizing the utilization of GPU parallel computing power.
3.1.4. GPG
Group Policy Gradient (GPG) (Chu et al., 2025a) eliminates the reliance on a reference model and avoids KL divergence constraints by directly optimizing the original RL objective, thereby simplifying the training process and reducing computational costs. Additionally, GPG introduces an Accurate Gradient Estimation (AGE) technique and a group reward normalization method, effectively addressing issues related to advantage function and gradient estimation bias. These innovations enhance training stability and performance.
Direct Optimization of the Original RL Objective.
Methods like GRPO typically rely on surrogate loss functions to approximate policy optimization. By contrast, GPG directly optimizes the original RL objective, avoiding biases that surrogate loss functions might introduce. Its core optimization objective is formulated as follows:
(11) $J_{\mathrm{GPG}}(\theta)=\mathbb{E}_{(q,a)\sim D,\{o_{i}\}_{i=1}^{G}}\left[\frac{1}{\sum_{i=1}^{G}\|o_{i}\|}\sum_{i=1}^{G}\sum_{t=1}^{\|o_{i}\|}\left(-\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})A_{i,t}\right)\right],$
where $A_{i,t}$ is the normalized intra-group advantage function.
Elimination of the Reference Model
As shown in Equation 5, GRPO stabilizes training by using a reference model and KL divergence constraints, whereas in Equation 11 GPG completely removes these components. This results in a more concise loss function for GPG.
Addressing Advantage Function and Gradient Estimation Bias
GPG proposes an Accurate Gradient Estimation (AGE) technique, which effectively resolves the issue of reward bias, formulated as:
(12) $\hat{g}=\frac{\sum_{i=M+1}^{B}g_{i}}{B-M}=g\cdot\alpha,\quad\alpha=\frac{B}{B-M},$
where $M$ represents the number of invalid samples (i.e., samples with zero gradients), and $\alpha$ is a dynamic adjustment factor used to correct gradient estimation bias.
3.2. Value-Model-Based
Value-model-based methods, exemplified by PPO,excel by enabling precise, step-wise credit assignment—an especially crucial feature for complex reasoning tasks, where minor errors at individual steps can propagate into catastrophic failures. In contrast to the advantage estimates from Monte Carlo methods in value-model-free approaches, a well-trained value model not only provides lower-variance token-level value estimates for enhanced training stability, but also exhibits inherent generalization capabilities that improve online sample utilization efficiency(Yuan et al., 2025a). While training value models for long CoT reasoning remains challenging, recent studies have introduced innovative optimization techniques that significantly improve both stability and overall performance.
3.2.1. ORZ
Open-Reasoner-Zero (ORZ) (Hu et al., 2025) demonstrates through extensive experiments that by using the vanilla PPO algorithm combined with GAE ($\lambda=1$, $\gamma=1$) and a simple rule-based reward function, significant improvements in response length and benchmark performance can be achieved without any KL regularization. Additionally, the authors highlight the importance of increasing both the quantity and diversity of training data for enhancing model performance. Compared to DeepSeek-R1-Zero, ORZ achieves comparable performance with only one-tenth of the training steps.
@@ 3.2.2. VC-PPO
To enable traditional reinforcement learning algorithms like PPO to handle long CoT tasks, Value-Calibrated PPO (VC-PPO) (Yuan et al., 2025b) introduces optimizations in two key areas: value initialization bias and Decoupled-GAE.
@@
The value function $V(s)$ in PPO is typically initialized using a reward model during the early stages of training. However, this initialization method can lead to significant bias, especially in long-sequence tasks. In the initial phase, since the reward model tends to underestimate the scores of earlier tokens, the advantage function $A_{t}$ exhibits a positive bias when estimating the contributions of these tokens. This bias encourages the policy to favor shorter outputs, thereby undermining the model’s ability to generate long CoT sequences. To address this issue, VC-PPO proposes Value Pretraining. By offline training the value model under a fixed policy and using GAE with $\lambda=1.0$ to ensure unbiased returns, the value model achieves a more accurate estimation of expected rewards.
@@
3.2.3. VAPO
Value-model-based Augmented Proximal Policy Optimization (VAPO) (Yuan et al., 2025a) enhances the training stability of PPO in long-chain CoT reasoning tasks. It adopts several tricks from DAPO (Yu et al., 2025b) and VC-PPO (Yuan et al., 2025b), such as the Clip-Higher Strategy, Token-Level Policy Gradient Loss, and Decoupled-GAE, while introducing Length-Adaptive GAE, Positive Example LM Loss, and Group-Sampling to achieve balanced optimization across varying sequence lengths and mitigate the issue of sparse rewards.
Length-Adaptive GAE
Building on the Decoupled-GAE method in VC-PPO, VAPO introduces a Length-Adaptive GAE approach that dynamically adjusts the $\lambda$ parameter to balance the bias-variance tradeoff for sequences of varying lengths. The specific formula is as follows:
(14) $\lambda_{\text{policy}}=1-\frac{1}{\alpha l},$
where $l$ denotes the sequence length and $\alpha$ is a hyperparameter that controls the overall bias-variance trade-off. This formula ensures a more uniform distribution of the TD-error coefficients, preventing the excessive decay of reward signals in long sequences.
Positive Example LM Loss
VAPO incorporates the Positive Example LM Loss, which leverages the negative log-likelihood (NLL) of the correct answers to improve the utilization efficiency of positive samples. The corresponding formula is:
(15) $\mathcal{L}_{\text{NLL}}(\theta)=-\frac{1}{\sum_{o_{i}\in T}\|o_{i}\|}\sum_{o_{i}\in T}\sum_{t=1}^{\|o_{i}\|}\log\pi_{\theta}(a_{t}\mid s_{t}),$
where $T$ denotes the set of correct answers, and $\|o_{i}\|$ represents the length of the sequence $o_{i}$.
Group-Sampling
VAPO employs Group-Sampling to extract discriminative positive and negative samples within the same prompt. It highlights that, given a fixed computational budget, reducing the number of distinct prompts per batch and reallocating resources for repeated generations improves performance. This is attributed to the introduction of richer contrastive signals, which enhances the learning capability of the policy model.
@@ 4. RL-Based MLLM Reasoning
KIMI K1.5 (Team et al., 2025a) KIMI K1.5 SFT&RL OPMD Univ. Reason T
@@
Visual-RFT (Liu et al., 2025f) Qwen2-VL-2B RL GRPO Detection&CLS T
VisualThinker-R1-Zero (Zhou et al., 2025a) Qwen2-VL-2B RL GRPO Spatial Reason T
@@
Vision-R1 (Zhan et al., 2025) Qwen2.5-VL-7B, Griffon-G-7B (Zhan et al., 2024) SFT&RL GRPO Detection T
Seg-Zero (Liu et al., 2025e) Qwen2.5-VL-3B RL GRPO Segmentation T
@@
MM-Eureka (Meng et al., 2025) InternVL2.5-Instruct-8/38B RL RLOO Math T
@@
R1-VL (Zhang et al., 2025a) Qwen2.5-VL-7B SFT&RL StepGRPO Univ. Reason T&I Desc
@@
OThink-MR1 (Liu et al., 2025g) Qwen2-VL-Instruct-2/7B RL GRPO-D Univ. Reason T
OpenVLThinker (Deng et al., 2025a) Qwen2.5-VL-Instruct-7B SFT&RL GRPO Math T
MetaSpatial (Pan and Liu, 2025) Qwen-VL-3B/7B RL GRPO Univ. Reason T
UI-R1 (Lu et al., 2025) Qwen2.5-VL-3B RL GRPO GUI T
@@
Q-Insight (Li et al., 2025d) Qwen-2.5-VL-7B RL GRPO Image Quality T&I Desc
Kimi-VL (Team et al., 2025b) Kimi-VL SFT&RL OPMD Univ. Reason T
ThinkLite-VL (Wang et al., 2025e) Qwen2.5-VL-Instruct-7B RL GRPO Univ. Reason T
Perception-R1 (Yu et al., 2025a) Qwen2-VL-Instruct-2B RL GRPO Univ. Reason T
VLAA-Thinker (Chen et al., 2025b) Qwen2-VL-2/7B, Qwen2.5-VL-3/7B RL GRPO Univ. Reason T
VL-Rethinker (Wang et al., 2025c) Qwen2.5-VL-7/72B RL GRPO-SSR Univ. Reason T
VLM-R1 (Shen et al., 2025) Qwen2.5-VL-Instruct-3/7/32B RL GRPO Univ. Reason T
GUI-R1 (Xia and Luo, 2025) Qwen2.5-VL-3/7B SFT&RL GRPO GUI T
InfiGUI-R1 (Liu et al., 2025c) Qwen2.5-VL-3B-Instruct SFT&RL RLOO GUI T
NoisyRollout (Liu et al., 2025d) Qwen2.5-VL-7B-Instruct RL GRPO Math T
R1-SGG (Chen et al., 2025c) Qwen2-VL-2/7B SFT&RL GRPO Generation T
Relation-R1 (Li et al., 2025a) Qwen2.5-VL-3B SFT&RL GRPO Spatial Reason T
Skywork R1V2 (Wei et al., 2025) InternViT-6B, QwQ-32B MPO&RL GRPO-SSB Univ. Reason T
FAST (Xiao et al., 2025) Qwen2.5-VL-3/7B RL Fast-GRPO Univ. Reason T
ChestX-Reasoner (Fan et al., 2025) Qwen2-VL-7B SFT&RL GRPO Medical T
@@
R1-Zero-VSI (Liao et al., 2025) Qwen2-VL-2/72B RL GRPO Spatial Reason T&I Desc
Spatial-R1 (Ouyang, 2025) Qwen2.5-VL-7B-Instruct RL GRPO Spatial Reason T
Embodied-R (Zhao et al., 2025a) Qwen2.5-3B-Instruct Qwen2.5-VL-72B-Instruct RL GRPO Spatial Reason T
TinyLLaVA-Video-R1 (Zhang et al., 2025b) Qwen2.5-3B&SigLIP RL GRPO ST-Perception T
VideoChat-R1 (Li et al., 2025c) Qwen2/2.5-VL-7B RL GRPO ST-Perception T
@@
R1-AQA (Li et al., 2025b) Qwen2-Audio-7B-Instruct RL GRPO Audio QA T
SARI (Wen et al., 2025) Qwen2-Audio-7B-Instruct Qwen2.5-Omni SFT&RL GRPO Audio QA T